# SCILLA: SurfaCe Implicit Learning for Large Urban Area, a volumetric hybrid solution

### v2

@@ 3.4 Optimization
3.4.1 Photometric loss.
We use the standard $L_{1}$ loss to minimize the pixel-wise color difference between the rendered image $\hat{C}$, and the ground truth image $C$, as well as DSSIM [33] loss on color patches.
3.4.2 Monocular cues.
As we target to reconstruct outdoor datasets, we add an additional sky loss to our model. Similar to StreetSurf [11], we model the sky color with an auxiliary MLP conditioned on the ray direction and use a loss to constrain the opacity of sky pixels to zero. The segmentation mask is obtained with an off-the-shelf semantic segmentation network [6].
@@
$\mathcal{L}_{\hat{N}}=\left\|\frac{\nabla f(x_{N})}{\left\|\nabla f(x_{N})\right\|_{2}}-\bar{N}(r)\right\|+\left\|1-\left(\frac{\nabla f(x_{N})}{\left\|\nabla f(x_{N})\right\|_{2}}\right)^{\top}\bar{N}(r)\right\|.$ (8)
3.4.3 Signed distance field regularization.
Similar to [34], to constrain our model to respect the eikonal equation [10] and obtain a more accurate surface reconstruction, we use the eikonal regularization defined as:
@@
According to NeuS formulation of the alpha compositing equation (Eq. 3), the network converges properly when the standard deviation $1/s$ of $\phi_{s}$ approaches zero. In complex driving scenarios and very large scenes, the parameter $s$ doesn’t always reach a satisfactory level of convergence. Unlike StreetSurf which manually fixes the parameter $s$ to increment linearly, we introduce a regularization term on $s$: $\mathcal{L}_{\text{s}}=1/(s+\epsilon)$.
@@ 4 Experiments
We evaluate the performance of SCILLA on multiple driving scenes and provide qualitative and quantitative comparisons compared to a relevant SoTA method.
4.0.1 Implementation details.
@@
4.0.2 Datasets.
We evaluate our method on four public autonomous driving datasets: KITTI-360 [18], nuScenes [3], Waymo Open Dataset [29] and Pandaset [38]. We conduct our experiments on four scenes from each one of these datasets. Sequences were selected to have large differences between them. All sequences are static, except for Pandaset where dynamic vehicles are masked during training. Complete details about the selected sequences can be found in our supplementary.
4.0.3 Baseline.
We benchmark our method against StreetSurf [11], which is the current SoTA method for driving scenes surface reconstruction. StreetSurf relies on hierarchical space partitioning and uses 3 distinct fields to model close range, far range and sky areas. It employs both 3D and 4D hash grid positional encoding as well as occupancy grids for efficient ray sampling. We used their official open-source implementation and evaluated their method without LiDAR supervision for a fair comparison. Trajectories are not optimized, relying only on monocular cues (depths and normals) for more challenging setups. All experiments were performed on the same computational setup for fair efficiency comparisons.
4.0.4 Evaluation metrics.
@@ 4.1 Results
(c) Waymo [29]
(d) nuScenes [3]
Figure 5: Mean cumulative delta error for both SCILLA and StreetSurf computed across the four sequences from each dataset as detailed in table 1. Curves indicate the percentage of ground-truth points having an error distance to the closest predicted mesh triangle which is lower than a given value. Light contours represent the standard deviation for each method.
@@
StreetSurf [11] 0.14 0.50 0.09 0.71 0.10 0.67 0.10 0.66 2.33 0.10 0.47 0.40 0.36 0.30 0.28 0.31
SCILLA (ours) 0.13 0.56 0.11 0.70 0.11 0.66 0.13 0.72 0.49 0.23 0.28 0.46 0.33 0.29 0.35 0.34
Waymo Open Dataset  [29] nuScenes [3]
Seq. 10061 Seq. 13196 Seq. 14869 Seq. 102751 Seq. 0034 Seq. 0071 Seq. 0664 Seq. 0916
@@
StreetSurf [11] 0.25 0.42 0.35 0.53 0.26 0.34 0.33 0.23 0.91 0.27 0.78 0.47 0.67 0.50 0.65 0.28
SCILLA (ours) 0.20 0.43 0.22 0.48 0.15 0.47 0.21 0.29 0.42 0.20 0.22 0.59 0.40 0.40 0.22 0.54
4.1.1 Quantitative analysis.
Table 2: Training analysis.
Iteration time GPU memory Params size
(s) (Gb) (MiB)
StreetSurf [11] 0.2 19 92.59
SCILLA (ours) 0.1 8 27
[…]
Efficiency. We report in Tab. 4.1.1 the computational performance of each method. Because SCILLA relies on a single 3D hierarchical hash grid in comparison to the two grids (3D and 4D) used in StreetSurf [11], our method is much more efficient by being designed with a third of the parameters of StreetSurf and taking half of the GPU memory. Furthermore, for a single scene, SCILLA requires half the training time compared to StreetSurf, on average.
KITTI-360 [18]
RGB Refer to caption Refer to caption Refer to caption
SCILLA (ours) Refer to caption Refer to caption Refer to caption
StreetSurf [11] Refer to caption Refer to caption Refer to caption
(a) Seq. 30 (b) Seq. 31 (c) Seq. 36
Figure 6: Qualitative experiments results on three sequences of KITTI-360. We report the RGB ground-truth image, from a different point-of-view than the RGB we compare our generated SDF mesh to StreetSurf’s mesh.
4.1.2 Qualitative analysis.
The quantitative evaluation presented in the previous section (Sec. 4.1.1) is computed from the cumulated ground-truth LiDAR which are sparse and do not cover some regions of the scene. To complete our evaluation, we present qualitative results in Figs.  6, 7, and 9. Results demonstrate that SCILLA reconstructs higher-quality surfaces compared to StreetSurf. We observe that our method recovers more accurate and realistic scene details at many regions such as buildings, cars, poles, and trees as it is highlighted by the red squares shown on each image. Our analysis supports the hypothesis that SteertSurf’s disentanglement of the close and far range based on the assumption that driving scenes are long and narrow is not adapted to the majority of driving sequences. In addition, StreetSurf’s road-surface initialization fails in sequences where the road and sidewalks are not at the same height or at non-flat roads such as downhills (see Fig. 6 – Seq. 30).
Pandaset [38]
RGB Refer to caption Refer to caption Refer to caption Refer to caption
StreetSurf [11] Refer to caption Refer to caption Refer to caption Refer to caption
(a) Seq. 23 (b) Seq. 37 (c) Seq. 42 (d) Seq. 43
Figure 7: Qualitative experiments results on four sequences of Pandaset. We report the masked RGB ground-truth image. In a different point-of-view than the GT we compare our generated SDF mesh to StreetSurf’s mesh.
nuScenes [3] Waymo Open Dataset [29]
RGB Refer to caption Refer to caption Refer to caption Refer to caption
StreetSurf [11] Refer to caption Refer to caption Refer to caption Refer to caption
(a) Seq. 71 (b) Seq. 916 (c) Seq. 10061 (d) Seq. 102751
Figure 8: Qualitative experiments on both nuScenes [3] (Left), and Waymo [29] (Right). We report the RGB ground-truth image. In a different point-of-view than the GT we compare our generated SDF mesh to StreetSurf’s mesh.
4.2 Ablation study
In this section, we ablate the effect of SCILLA’s hybrid volume rendering strategy and show quantitative and qualitative results in Tab. 3. The results demonstrate that transitioning from the volumetric to the surface representation is crucial to accurately reconstruct the road and various scene aspects such as signs and trees. In addition, we report our method with standard accumulated normal supervision (Sec. 3.4.2), and without the network regularization loss (Sec. 3.3.3).
[…]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GT RGB (a) w/o Hybrid (b) w/o $\mathcal{L}_{\hat{N}}$ (c) w/o $\mathcal{L}_{\text{s}}$ (d) SCILLA
[…]
Prec. 0.38 0.31 0.42 0.54
@@ 5 Discussion
5.0.1 Applications.
Thanks to the quality and completeness of SCILLA’s reconstructed meshes, we can use modern MVS tools such as OpenMVS [4] to obtain a detailed and colorized representation of large urban scenes, as shown in Fig. 1. In addition, SCILLA is not restricted to driving scenarios, it can be applied to other types of scenes. We encourage the reader to refer to the supplementary materials for qualitative results of SCILLA on the Tanks&Temples dataset [16].
5.0.2 Limitations.
Although our method can achieve highly detailed surface reconstruction in driving scenarios, it fails to reconstruct fine details in very wide and open sequences. We recommend consulting the supplementary materials for the failure cases of SCILLA.
@@ 6 Conclusion
Appendix A Additional implementation details
The overall loss we use to optimize SCILLA is defined as follows:
$\mathcal{L}=\mathcal{L}_{\text{rgb}}+\lambda_{1}\mathcal{L}_{\text{dssim}}+\lambda_{2}\mathcal{L}_{\hat{N}}+\lambda_{3}\mathcal{L}_{\text{eik}}+\lambda_{4}\mathcal{L}_{\text{sky}}$ (10)

### v3

@@ 3.4 Optimization
Photometric loss. We use the standard $L_{1}$ loss to minimize the pixel-wise color difference between the rendered image $\hat{C}$, and the ground truth image $C$, as well as DSSIM [36] loss on color patches.
Monocular cues. As we target to reconstruct outdoor datasets, we add an additional sky loss to our model. Similar to StreetSurf [11], we model the sky color with an auxiliary MLP conditioned on the ray direction and use a loss to constrain the opacity of sky pixels to zero. The segmentation mask is obtained with an off-the-shelf semantic segmentation network [6].
@@
$\mathcal{L}_{\hat{N}}=\left\|\frac{\nabla f(x_{N})}{\left\|\nabla f(x_{N})\right\|_{2}}-\bar{N}(r)\right\|+\\ \left\|1-\left(\frac{\nabla f(x_{N})}{\left\|\nabla f(x_{N})\right\|_{2}}\right)^{\top}\bar{N}(r)\right\|.$ (7)
Signed distance field regularization. Similar to [37], to constrain our model to respect the eikonal equation [10] we use the eikonal regularization defined as:
@@
According to NeuS formulation of the alpha compositing equation (Eq. 3), the network converges properly when the standard deviation $1/s$ of $\phi_{s}$ approaches zero. In complex driving scenarios and very large scenes, the parameter $s$ does not always reach a satisfactory level of convergence. Unlike StreetSurf which manually fixes the parameter $s$ to increment linearly, we introduce a regularization term on $s$: $\mathcal{L}_{\text{s}}=1/(s+\epsilon)$.
KITTI-360 [20] Pandaset [41] Waymo [32] nuScenes [3]
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Mean cumulative delta error for both SCILLA and StreetSurf computed across the four sequences from each dataset as detailed in Tab. 1. Curves indicate the percentage of ground-truth points having an error distance to the closest predicted mesh triangle which is lower than a given value. Light contours represent the standard deviation for each method.
KITTI-360 [20] Pandaset [41]
Seq. 30 Seq. 31 Seq. 35 Seq. 36 Seq. 23 Seq. 37 Seq. 42 Seq. 43
[…]
StreetSurf [11] 0.19 0.50 0.15 0.71 0.16 0.67 0.15 0.66 2.35 0.10 0.48 0.4 0.40 0.30 0.28 0.31
SCILLA (ours) 0.18 0.56 0.15 0.71 0.15 0.66 0.16 0.72 0.51 0.23 0.29 0.46 0.36 0.29 0.35 0.34
COLMAP [30] 0.15 0.70 0.12 0.81 0.12 0.78 0.15 0.74 0.71 0.41 1.0 0.52 0.78 0.53 0.28 0.31
OpenMVS [4] 0.13 0.77 0.12 0.80 0.11 0.82 0.13 0.80 0.65 0.56 1.02 0.53 0.74 0.58 0.28 0.31
Waymo [32] nuScenes [3]
Seq. 10061 Seq. 13196 Seq. 14869 Seq. 102751 Seq. 0034 Seq. 0071 Seq. 0664 Seq. 0916
[…]
StreetSurf [11] 0.26 0.42 0.37 0.53 0.25 0.35 0.30 0.23 0.93 0.27 0.80 0.47 0.70 0.50 0.67 0.28
SCILLA (ours) 0.23 0.43 0.25 0.48 0.20 0.47 0.24 0.30 0.46 0.20 0.25 0.59 0.44 0.40 0.25 0.54
COLMAP [30] 0.27 0.56 0.78 0.53 0.19 0.67 0.27 0.58 0.67 0.48 0.85 0.67 0.80 0.62 0.80 0.68
OpenMVS [4] 0.3 0.57 0.76 0.53 0.21 0.67 0.24 0.63 0.62 0.51 0.86 0.69 0.76 0.70 0.75 0.69
[…]
@@ 4 Experiments
We evaluate the performance of SCILLA on multiple driving scenes and provide qualitative and quantitative comparisons compared to relevant methods.
4.1 Implementation details
@@
4.2 Datasets
We evaluate our method on four public driving datasets: KITTI-360 [20], nuScenes [3], Waymo Open Dataset [32] and Pandaset [41]. We conduct our experiments on four scenes from each dataset. Sequences were selected to have large differences between them. All sequences are static, except for Pandaset where dynamic vehicles are masked during training. Complete details about the selected sequences can be found in our supplementary.
4.3 Baseline
StreetSurf [11] is the current SoTA for surface reconstruction in driving scenes, using road-based surface initialization tailored for such sequences. It employs hierarchical space partitioning with 3 distinct fields to model close-range, far-range, and sky areas, along with 3D and 4D hash grid positional encoding and occupancy grids for efficient ray sampling. We used the official open-source implementation, evaluated with monocular depth supervision, and deactivated trajectory optimization for a fair comparison. All experiments were conducted on the same computational setup. We also extensively tested Neuralangelo [19] on the selected driving sequences with various training strategies and hyperparameters tuning but did not achieve satisfactory results. Neuralangelo’s need for a bounded scene to initialize the SDF in a sphere shape makes it unsuitable for the open, unbounded nature of driving sequences. Further details can be found in our supplementary materials. For a comprehensive comparison, we benchmark SCILLA’s mesh reconstruction against two classic MVS methods: Delaunay triangulation in COLMAP [30] and mesh reconstructionn [15] from OpenMVS [4], both using the denser point cloud from COLMAP.
4.4 Evaluation metrics
@@ 4.5 Results
Quantitative analysis.
[…]
Efficiency. We report in Tab. 2 the computational performance of each method. SCILLA uses a single 3D hierarchical hash grid, unlike StreetSurf’s two grids (3D and 4D) [11], making it more efficient with one-third of the parameters and half the GPU memory during training. Additionally, SCILLA requires half the training time per scene compared to StreetSurf.
Total train time GPU memory Params size
(min) (Gb) (MiB)
StreetSurf [11] 40 19 92.59
SCILLA (ours) 20 8 27
COLMAP [30] 30 $<$1
OpenMVS [4] 30 $<$1
Table 2: Training analysis.
Qualitative analysis.
The quantitative evaluation presented in the previous section (Sec. 4.5) is computed from the accumulated ground-truth LiDAR which is sparse and may miss some scene regions. To complete our evaluation, we present qualitative results in Fig. 8. Results demonstrate that SCILLA reconstructs higher-quality surfaces compared to StreetSurf and MVS methods. MVS-generated meshes have noisy normals and are incomplete. We observe that our method recovers more accurate and realistic scene details in many regions such as buildings, cars, poles, and trees. Our analysis supports the hypothesis that SteertSurf’s disentanglement of the close and far range based on the assumption that driving scenes are long and narrow is not adapted to the majority of driving sequences. In addition, StreetSurf’s road-surface initialization fails in sequences where the road and sidewalks are not at the same height or at non-flat roads such as downhills (see Fig. 8 – Seq. 30).
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
Seq. 30 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seq. 31 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seq. 36 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
Seq. 23 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seq. 42 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seq. 43 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
Seq. 71 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seq. 916 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(c) nuScenes [3]
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
Seq. 71 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seq. 916 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(d) Waymo Open Dataset [32]
Figure 5: Qualitative experiments results on 8(a) KITTI-360s [20], 8(b) Pandaset [41], 8(c) nuScenes [3] and 8(d) Waymo Open Dataset [32]. RGB ground-truth images shown for reference of the scene layout. Complex scene geometries reconstructed in the mesh are highlighted in red squares. We compare our generated SDF mesh to COLMAP, OpenMVS and StreetSurf [11] meshes.
4.6 Ablation study
Convergence of density field compared to SDF field. We report in Fig. 6(a) the Chamfer Distance between LiDAR and predicted point clouds, for Seq. 31 from KITTI-360 and Seq. 916 from nuScenes, comparing two different NeRF: one solely based on density volumetric field and another one solely based on SDF at various training steps (without normal supervision, as we cannot supervise normal in case of pure density volumetric field). The results support the claim that density field converges faster compared to the SDF field.
SCILLA’s components. In this section, we ablate SCILLA’s SDF progressive initialization strategy, with quantitative results reported in Tab. 3. The results confirm that transitioning from volumetric to surface representation is essential for accurately initializing the SDF field and efficiently reconstructing driving sequences. We also report results with standard accumulated normal supervision and without network regularization loss.
@@ 4.6 Ablation study
Table 3: Ablations study: we deactivate some SCILLA ’s key components. (a) without progressive initialization (b) with the accumulated normal for $\mathcal{L}_{\hat{N}}$ (c) without $\mathcal{L}_{\text{s}}$.
w/o SDF init w/o $\mathcal{L}_{\hat{N}}$ w/o $\mathcal{L}_{\text{s}}$ SCILLA
@@
KITTI-360. 31 0.35 0.18 0.19 0.27 0.24 0.15 0.15 0.71
Pandaset. 23 0.85 0.18 0.80 0.18 0.54 0.17 0.51 0.23
nuScenes. 0916 0.35 0.4 0.4 0.21 0.35 0.15 0.25 0.54
Waymo. 10061 0.3 0.36 0.22 0.32 0.34 0.17 0.23 0.43
@@ 5 Discussion
Applications. Thanks to the quality and completeness of SCILLA’s reconstructed meshes, we can use modern MVS tools such as OpenMVS [4] to obtain a detailed and colorized representation of large urban scenes, as shown in Fig. 1. In addition, SCILLA is not restricted to driving scenarios, it can be applied to other types of scenes. We encourage the reader to refer to the supplementary materials for qualitative results of SCILLA on the Tanks&Temples dataset [18].
Limitations. Although our method can achieve highly detailed surface reconstruction in driving scenarios, it fails to reconstruct fine details in very wide and open sequences. We recommend consulting the supplementary materials for the failure cases of SCILLA.
@@ 6 Conclusion
[…]
Supplementary Material
In this supplementary material, we provide additional implementation details, experiments with Neuralangelo [19], additional qualitative results, ablation study on our samples attribution strategy, applications, and failure cases of SCILLA.
7 SDF-gradient normalization
As introduced in Sec. 3.3 of the main paper, volumetric representations based on density estimation tend to quickly approximate saturated alpha values ($\alpha_{i}^{v}$ being either 00 or $1$), while alpha values computed from the SDF converge slower. Because $\alpha_{i}^{v}$ and $\alpha_{i}^{f}$ are composed jointly during the hybrid stage, we noticed that surface representation compensates for this gap by predicting large SDF gradient values so that $\alpha_{i}^{f}$ aligns with $\alpha_{i}^{v}$. Indeed, from Eq. 3 of the main paper, a solution to saturate alpha towards either 1 or 0, is to predict $f(p_{i+1})\ll f(p_{i})$ or $f(p_{i+1})\gg f(p_{i})$, respectively. During the forward pass, $f(p_{i})$ and $f(p_{i+1})$ are not directly predicted by the model but rather computed using:
with $\theta$ being the angle between the direction of $\nabla f(x_{i})$ and the ray direction $d$. Practically, we do not formally compute $\theta$ but we rather approximate its cosine with $\cos({\theta})=\nabla f(x_{i})\cdot d$, assuming both vectors are unit norm. By predicting $\left\|\nabla f(x_{i})\right\|_{2}\gg 1$, $\alpha_{i}^{f}$ can be easily saturated and follow the distribution of $\alpha_{i}^{v}$ without learning a proper signed distance field. To address this, we simply normalize the SDF gradient before the cosine computation to prevent the gradient from compensating the alpha distribution difference. It is important to notice that even if the eikonal loss used for training (see main paper) is supposed to encourage the network to model a signed distance function with spatial derivative of unitary norm, we found that numerically normalizing the gradient during our hybrid stage is essential to avoid divergence in early training iterations.
Neuralangelo [19] SCILLA (ours)
Refer to caption Refer to caption
Refer to caption Refer to caption
(a) Seq 30. from KITTI-360 [20]
Neuralangelo [19] SCILLA (ours)
Refer to caption Refer to caption
Refer to caption Refer to caption
(b) Seq. 31 from KITTI-360 [20]
Full seq. Segmented sequences
Refer to caption Refer to caption
(c) Failure of Neuralangelo on Seq 10061. from Waymo Open Dataset [32]
Figure 7: Qualitative experiments results on Seq. 30 7(a) and Seq. 31 7(b) from KITTI-360s [20]. Failure on Seq. 10061 7(c) Waymo Open Dataset [32]. We compare our generated SDF mesh to Neuralangelo [19] mesh for KITTI-360. We report the experiments on the full and segmented sequence for Waymo.
8 SDF Field Initialization: Neuralangelo
As detailed in the main paper, Neuralangelo [19] is designed for landmark reconstruction, relying on many overlapping images and a bounded region of interest to initialize the SDF with a spherical shape. In our initial tests, when applying Neuralangelo (using the official authors codebase  [^1] https://github.com/NVlabs/neuralangelo) to outdoor driving scenarios using the default settings for outdoor scenes, we observed that the initial part of the scene remained noisy and retained a spherical shape. Only the final part of the scene, which was consistently visible across all images in the sequence, was accurately reconstructed (see Fig. 7(a)). While Neuralangelo managed to reconstruct some parts of simpler scenes, such as in Seq. 31 from KITTI-360 (Fig.7(b)), it failed completely in more complex cases, particularly in sequences that are long, wide, or contain challenging structure like downhills (e.g., Seq. 30 in Fig.7(a)). Moreover, Neuralangelo was unable to reconstruct any part of the Waymo dataset, as shown in Fig. 7(c). Given these outcomes, we conducted extensive experiments with Neuralangelo using various training strategies:
  • Constraining the spherical shape: since Neuralangelo relies on a spherical initialization where everything outside the sphere is treated as background, we attempted to center the sphere on a smaller region, with a smaller radius. However, this approach resulted in very noisy reconstructions, likely due to insufficient overlapping images in that part of the scene (inside the sphere).
  • Segmenting the scene into multiple parts: acknowledging that driving sequences are typically long and wide, we divided the scenes into smaller segments to better fit the spherical initialization. This strategy, however, produced unsatisfactory results. Neuralangelo requires a large number of images (typically 300 for an average Tanks and Temples scene covering the same region of interest) while KITTI-360 sequences contain around 200 images (approximately 50 per camera) and after segmentation, each part of the scene had roughly 100 images. The method could only reconstruct the final segment of the scene that was visible in all images (see Fig. 7(c)).
  • Doubling the number of images: we further experimented with doubling the number of images used in the reconstruction. Unfortunately, this did not help the method to converge to a satisfactory reconstruction.
Despite various training strategies, Neuralangelo failed to produce reliable reconstructions for complex driving sequences. It requires a large number of overlapping images and a tightly bounded scene, making it unsuitable for unbounded driving sequences. Additionally, Neuralangelo’s training time is prohibitively long, taking up to 24 hours per scene. This, combined with its poor performance on driving sequences, makes it impractical for large-scale driving scene reconstruction tasks.
9 Additional implementation details
We use the poses provided by the datasets, except for Waymo, where we recompute the vehicle trajectory and sensor calibration with MOISST [13] due to inaccuracies in the provided data. The overall loss we use to optimize SCILLA is defined as follows:
$\mathcal{L}=\mathcal{L}_{\text{rgb}}+\lambda_{1}\mathcal{L}_{\text{dssim}}+\lambda_{2}\mathcal{L}_{\hat{N}}+\lambda_{3}\mathcal{L}_{\text{eik}}+\lambda_{4}\mathcal{L}_{\text{sky}}.$ (10)